Robert invests $8000 in a bank. Part is invested at 7% simple interest and part at 9% simple interest. In one year Robert earns $660 in interest. How much did he invest at each amount?
Answer by rfer(10425)

.07x+.09(8000-x)=660
.07x+720-.09x=660
-.02x=-60
x=$3000 at 7%
8000-3000=$5000 at 9%